caira12:

10y + 3x + 10 +x -2y = ?

Vocaloid:

start by combining the x terms 3x + x = ?

Vocaloid:

3 x's plus 1x gives 4 x's so 3x + x = 4x

Vocaloid:

using that same logic 10y - 2y = ?

caira12:

8y

Vocaloid:

good so putting everything together 10y + 3x + 10 +x -2y = 4x + 8y + 10 and since all the like terms are combined, that's all that can be done